OpenStudy (acxbox22):

a farmer has 120 yards of fencing to build around a rectangular orchard. let w be the width of the orchard. writean equation giving the area of the orchard. find the dimensions of the orchard when the area is 900 square yards.

OpenStudy (acxbox22):

@SolomonZelman , @jim_thompson5910

OpenStudy (acxbox22):

@jdoe0001

jimthompson5910 (jim_thompson5910):

Assuming he uses all of the fencing to enclose the entire orchard, this means the perimeter of the orchard is 120 yards P = 120

jimthompson5910 (jim_thompson5910):

The perimeter of a rectangle is P = 2*(L+W) plug in P = 120 120 = 2*(L+W) then divide both sides by 2 60 = L + W and then solve for L to get L = 60 - W

jimthompson5910 (jim_thompson5910):

The area of a rectangle is A = L*W Plug in the desired area (900) for A and L = 60 - W to get 900 = (60 - W)*W and I'll let you take over from here

OpenStudy (acxbox22):

thanks

jimthompson5910 (jim_thompson5910):

np